# DPOT: Auto-Regressive Denoising Operator Transformer for Large-Scale PDE Pre-Training

### v2

@@ C.2 Proof of the theorem
Theorem C.1 (Universal Approximation by DPOT).
Let $s,s^{\prime}>0$. Let $\mathbb{T}^{d}=[0,2\pi]^{d}$ be the d-dimensional torus. Let $\mathcal{G}:H^{s}(\mathbb{T}^{d};\mathbb{R}^{d_{in}})\to H^{s^{\prime}}(\mathbb{T}^{d};\mathbb{R}^{d_{out}})$ be a continuous operator between Sobolev spaces, and $K\subset H^{s}(\mathbb{T}^{d};\mathbb{R}^{d_{in}})$ be a compact subset. Then for any $\varepsilon>0$, there exists a DPOT $\mathcal{N}:H^{s}(\mathbb{T}^{d};\mathbb{R}^{d_{in}})\to H^{s^{\prime}}(\mathbb{T}^{d};\mathbb{R}^{d_{out}})$ satisfies:
$\sup_{v\in K}\|\mathcal{G}(v)-\mathcal{N}(v)\|_{L^{2}}\leq\varepsilon$ (30)
@@
We start the proof by defining the following operator:
$\mathcal{G}_{N}:H^{s}(\mathbb{T}^{d})\to H^{s^{\prime}}(\mathbb{T}^{d}),\ \ \mathcal{G}_{N}(v)=\mathcal{P}_{N}\mathcal{G}(\mathcal{P}_{N}(v))$ (31)
Intuitively, the defined operator $\mathcal{G}_{N}$ is an approximation to the original operator $\mathcal{G}$ by truncating the high-frequency part of input and output functions. Thanks to Theorem 5 in (Kovachki et al., 2021), we have that for $\forall\varepsilon>0$, there exists $N\geq 0$ that:
$\|\mathcal{G}(v)-\mathcal{G}_{N}(v)\|_{L^{2}}\leq\varepsilon,\ \ \forall v\in K$ (32)
Thus, the main objective of the following proof is to construct a DPOT $\mathcal{N}$ to approximate $\mathcal{G}_{N}$.
Next, we define the Fourier Conjugate of $\mathcal{G}_{N}$ as:
$\hat{\mathcal{G}}_{N}:\mathbb{C}^{\mathcal{K}_{N}}\to\mathbb{C}^{\mathcal{K}_{N}},\ \ \hat{\mathcal{G}}_{N}(\hat{v})=\mathcal{F}(\mathcal{G}_{N}(\Re(\mathcal{F}^{-1}(\hat{v})),$ (33)
and it is easy to verify that:
$\mathcal{G}_{N}=\mathcal{F}^{-1}\circ\hat{\mathcal{G}}_{N}\circ(\mathcal{F}\circ\mathcal{P}_{N})$ (34)
Given the decomposition above, we can construct a DPOT by separately approximating the following three parts $\mathcal{F}^{-1}$, $\hat{\mathcal{G}}_{N}$ and $\mathcal{F}\circ\mathcal{P}_{N}$. The subsequent proof will consist of three lemmas, each addressing one of the aforementioned components.
To start with, we introduce the following lemma to fit $\hat{\mathcal{G}}_{N}$ with several Fourier Attention Layers of DPOT:
Lemma C.2 (Approximating $\hat{\mathcal{G}}_{N}$).
For any Fourier Conjugate Operator $\hat{\mathcal{G}}_{N}:\mathbb{C}^{\mathcal{K}_{N}}\to\mathbb{C}^{\mathcal{K}_{N}}$ and for any $\varepsilon>0$, we can construct $l$ DPOT layers $\mathcal{L}_{l}\circ\mathcal{L}_{l-1}\circ\cdots\circ\mathcal{L}_{1}$ that satisfy:
$|\hat{\mathcal{G}}_{N}(\hat{v})(x)-(\mathcal{L}_{l}\circ\mathcal{L}_{l-1}\circ\cdots\circ\mathcal{L}_{1})(\hat{w})(x)|_{2}\leq\varepsilon,\ \ \forall v\in K,\forall x\in\mathbb{T}^{d}$ (35)
where $|\cdot|_{2}$ is the 2-norm in vector space $\mathbb{C}^{\mathcal{K}_{N}}$ and $\hat{w}:\mathbb{T}^{d}\to\mathbb{C}^{\mathcal{K}_{N}},x\mapsto\hat{v}$ is a constant function defined on $\mathbb{T}^{d}$.
Remark C.3.
The reason why we introduce a constant function $\hat{w}$ is that operators are mappings between function spaces, but $\hat{\mathcal{G}}_{N}$ is a mapping between vector spaces. By using $\hat{w}$, we can address the issue of different domains.
@@
According to the definition of Fourier attention layers:
$\mathcal{L}_{l}(v)(x)=v(x)+\mathcal{M}_{l}\circ\mathcal{F}^{-1}(K_{l}(k)(\mathcal{F}(\mathcal{P}_{N}v)(k)))$ (36)
As $\mathcal{M}_{l}(v)(x)=M_{l}(v(x))$ is a 2-layer MLP with non-linear activation, we can leverage the Universal Approximation Theorem of Neural Networks (Funahashi, 1989) to fit the mapping $\hat{\mathcal{G}}_{N}$. Therefore, if we can set
$\mathcal{F}^{-1}(K_{l}(k)(\mathcal{F}(\mathcal{P}_{N}v)(k)))\approx\mathcal{P}_{N}v,\forall l,$ (37)
then we will have:
$\mathcal{L}_{l}(v)(x)\approx v(x)+M_{l}(\mathcal{P}_{N}v(x))$ (38)
and the Universal Approximation Theorem will guarantee that $\mathcal{L}_{l}\circ\mathcal{L}_{l-1}\circ\cdots\circ\mathcal{L}_{1}$ can approximate $\hat{\mathcal{G}}_{N}$ to any desired precision.
To fulfill the conditions mentioned earlier, we simply need to set $K_{l}(k)$ to the identity. Considering that $K_{l}(k)$ is a 2-layer block attention MLP with GELU activation, we can assign large positive values to the bias term in the first layer and circumvent the non-linearity in activation, thereby achieving the desired identity map following Eq. (1.8) in (Berner et al., 2021).
For the second step, we present the lemma to fit $\mathcal{F}\circ\mathcal{P}_{N}$:
Lemma C.4 (Approximating $\mathcal{F}\circ\mathcal{P}_{N}$).
For any $\varepsilon>0$, there exists $l\geq 0$ and $\mathcal{L}_{l}\circ\mathcal{L}_{l-1}\circ\cdots\circ\mathcal{L}_{1}\circ\mathcal{R}$ that satisfies:
$|\mathcal{F}(\mathcal{P}_{N}v)-(\mathcal{L}_{l}\circ\mathcal{L}_{l-1}\circ\cdots\circ\mathcal{L}_{1}\circ\mathcal{R})(v)(x)|_{2}\leq\varepsilon,\ \ \forall v\in K,\forall x\in\mathbb{T}^{d}$ (39)
where $|\cdot|_{2}$ is the 2-norm in vector space.
Proof.
The proof will consist of 2 steps. In the first step, we will show that we can construct $v(x)\cdot\sin(\langle k,x\rangle)$ and $v(x)\cdot\cos(\langle k,x\rangle)$. In the second step, we will leverage the Fourier transform to integrate these functions over the domain to get Fourier coefficients.
Step 1. First, we will show that there exists $\mathcal{N}=\mathcal{L}_{k}\circ\mathcal{L}_{k-1}\circ\cdots\circ\mathcal{L}_{1}\circ\mathcal{R}$ that satisfies:
$\begin{cases}\|\mathcal{N}(v)_{1,k}-P_{N}v(x)\cdot\cos(\langle k,x\rangle)\|_{L^{2}}<\varepsilon\\ \|\mathcal{N}(v)_{2,k}-P_{N}v(x)\cdot\sin(\langle k,x\rangle)\|_{L^{2}}<\varepsilon\\ \end{cases}$ (40)
To construct $\mathcal{N}$, we first leverage the position embedding layer and define the following lifting operator $\mathcal{R}$:
$\mathcal{R}(v)=R(v(x))=\{v(x),\cos(\langle k,x\rangle),v(x),\sin(\langle k,x\rangle)\}_{k\in\mathcal{K}_{N}}$ (41)
Here, the trigonometric functions are embedded directly within the embedding layer of DPOT. This lifting operator elevates the range of the function from $\mathbb{R}^{d_{in}}$ to $\mathbb{R}^{d_{hid}}$, where $d_{hid}=|\mathcal{K}_{N}|\times(2d_{in}+2)$.
Next, to construct the objective function, we need to follow a similar approach to the proof strategy in the previous lemma C.2. We set the frequency feature MLP $K$ in $\mathcal{L}$ to identity, thereby considering $\mathcal{L}$ as an MLP layer. Then, by leveraging the Universal Approximation Theorem of neural networks, we can create multiple layers to satisfy:
$(\mathcal{L}_{l}\circ\cdots\circ\mathcal{L}_{1})(\{v,\cos(\langle k,x\rangle),v,\sin(\langle k,x\rangle)\}_{k\in\mathcal{K}_{N}})\approx\{v\cdot\cos(\langle k,x\rangle),v\cdot\sin(\langle k,x\rangle)\}_{k\in\mathcal{K}_{N}},\ \ \forall v\in K,$ (42)
and achieve the desired precision by adjusting the width and depth of the layers.
Step 2. Second, we will utilize the Fourier Transform to obtain the Fourier coefficients.
By the definition of Fourier transform in Eq. 22:
$\mathcal{F}(v)(0)=\int_{\mathbb{T}^{d}}v(x)e^{\langle 0,x\rangle}\text{d}x=\int_{\mathbb{T}^{d}}v(x)\text{d}x$ (43)
Using this property, we can integrate the previously constructed functions, thus obtaining:
$\mathcal{F}(v\cdot\cos(\langle k,x\rangle))(0)=\int_{\mathbb{T}^{d}}v(x)\cos(\langle k,x\rangle)\text{d}x=\text{Re}(\hat{v}_{k})$ (44)
$\mathcal{F}(v\cdot\sin(\langle k,x\rangle))(0)=\int_{\mathbb{T}^{d}}v(x)\sin(\langle k,x\rangle)\text{d}x=\text{Im}(\hat{v}_{k})$ (45)
Thus, in the following layer $\mathcal{L}_{l+1}$, we can set:
$\begin{cases}K_{l+1}(k)=0&\text{if }k=0\\ K_{l+1}(k)=-Id&\text{if }k\neq 0\end{cases},$ (46)
and we will have:
$\mathcal{F}^{-1}(K_{l+1}(k)\cdot\mathcal{F}(\{v\cdot\cos(\langle k,x\rangle),v\cdot\sin(\langle k,x\rangle)\}_{k\in\mathcal{K}_{N}}))=\{\text{Re}(\hat{v}_{k})-v\cdot\cos(\langle k,x\rangle),\text{Im}(\hat{v}_{k})-v\cdot\sin(\langle k,x\rangle)\}_{k\in\mathcal{K}_{n}}$ (47)
Finally, we set the MLP layers $\mathcal{M}_{l+1}$ in $\mathcal{L}_{l+1}$ to an (approximate) identity, so that we can obtain:
$\mathcal{L}_{l+1}(\{v\cdot\cos(\langle k,x\rangle),v\cdot\sin(\langle k,x\rangle)\}_{k\in\mathcal{K}_{N}})=\{\text{Re}(\hat{v}_{k}),\text{Im}(\hat{v}_{k})\}_{k\in\mathcal{K}_{n}}=\mathcal{F}(\mathcal{P}_{N}v),$ (48)
which is the desired result.
For the last step, we will present the following lemma to approximate $\mathcal{F}^{-1}$:
Lemma C.5 (Approximating $\mathcal{F}^{-1}$).
For any $\varepsilon>0$, there exists $l\geq 0$ and $\mathcal{Q}\circ\mathcal{L}_{l}\circ\mathcal{L}_{l-1}\circ\cdots\circ\mathcal{L}_{1}$ that satisfies:
$\|\mathcal{F}^{-1}(\hat{v})-(\mathcal{Q}\circ\mathcal{L}_{l}\circ\mathcal{L}_{l-1}\circ\cdots\circ\mathcal{L}_{1})(v)\|_{L^{2}}\leq\varepsilon,\ \ \forall v\in K,$ (49)
where $\hat{v}\in\mathbb{C}^{\mathcal{K}_{N}}$ is the truncated Fourier coefficient of the function $v$.
Proof.
The proof of this lemma is similar to that of lemma C.4. We will first construct $\text{Re}(\hat{v}_{k})\cdot\cos(\langle k,x\rangle)$, $\text{Im}(\hat{v}_{k})\cdot\sin(\langle k,x\rangle)$ using $\mathcal{L}_{l}\circ\cdots\circ\mathcal{L}_{1}$ by repeating the proof in lemma C.4. Then, we can construct the following projection operator $\mathcal{Q}$ to recover the original function:
$\mathcal{Q}(\{\text{Re}(\hat{v}_{k})\cdot\cos(\langle k,x\rangle),\text{Im}(\hat{v}_{k})\cdot\sin(\langle k,x\rangle)\}_{k\in\mathcal{K}_{N}})=\sum_{k\in\mathcal{K}_{N}}\text{Re}(\hat{v}_{k})\cdot\cos(\langle k,x\rangle)-\text{Im}(\hat{v}_{k})\cdot\sin(\langle k,x\rangle)=v(x)$ (50)
Therefore, our proof is complete. ∎
The three lemmas above each demonstrate that we can approximate different parts of $\mathcal{G}_{N}$ with DPOT to any desired precision. Consequently, by combining these three theorems, we can approximate $\mathcal{G}_{N}$ to arbitrary precision. In this way, we have established the Uniform Approximation Theorem for DPOT.

### v4

@@ C.2 Proof of the theorem
In the following proof, we will demonstrate that there exists a DPOT model $\mathcal{N}$ that can approximate any operator $\mathcal{G}$ on a $d$-dimensional torus $\mathbb{T}^{d}$ with mesh $\mathcal{M}$. That is, for any $v\in L_{2}(\mathbb{T}^{d})$ and $\varepsilon>0$, we have:
$\|\mathcal{N}(\{v_{\mathbf{j}}\})-\mathcal{G}(\{v_{\mathbf{j}}\})\|_{\infty}<\varepsilon$ (33)
where $\{v_{\mathbf{j}}\}=\{v(x_{\mathbf{j}})\}_{j\in\mathcal{I}}$ is the value of $v$ on the mesh $\mathcal{M}\subset\mathbb{T}^{d}$.
The proof will consist of 3 steps. We will first construct a proxy target that is equivariant to input order, then construct a Sumformer to approximate the proxy leveraging Theorem C.7. Finally, we will show that there exists a corresponding DPOT that can fit the constructed Sumformer.
Step 1. Construction of proxy target.
As the first preparatory step, we will first assign a integer identifier to each point in the grid, thereby treating it as a 1 dimensional sequence. To be specific, let $N=n^{d}$ and we can map any indices $\mathbf{j}\in\mathcal{I}$ to an integer by:
[…]
In the subsequent proof, we will use new integer indices $j\in[0,N]$ to replace the original vector indices $\mathbf{j}\in\mathcal{I}$.
Next, in order to leverage Theorem C.7, we will construct an equivariant function by concatenating a permutation into the input. To be specific, let $\mathcal{\tilde{X}}=X\oplus\mathbb{N}$ and $U=\{[(x_{1},\pi_{1}),\cdots,(x_{N},\pi_{N})]|x_{j}\in\mathcal{X},\pi\in[N]\}\subset\mathcal{\tilde{X}}$, where $[N]$ is a permutation of $\{1,2,...,N\}$ the proxy target can be defined as:
[…]
where $\pi^{-1}$ is the inverse of $\pi$ in the permutation space.
Now, we will show the equivariance of $f$ by the following lemma:
Lemma C.8 (equivariance of $f$).
For any permutation $\sigma\in[N]$ and any input sequence $\{(x_{j},\pi_{j})\}\in U$, the proxy target $f$ satisfies:
$f([(x_{\sigma(1)},\pi_{\sigma(1)}),\cdots,(x_{\sigma(n)},\pi_{\sigma(N)})])=\sigma^{-1}\circ f([(x_{1},\pi_{1}),\cdots,(x_{N},\pi_{N})])$ (36)
@@
Let $\pi^{\prime}=\pi\circ\sigma$, it can be shown that:
[…]
Step 2. construct a Sumformer.
Given the equivariant sequence-to-sequence function $f$, we can now leverage Theorem C.7 to construct a Sumformer $\mathcal{S}$ satisfies:
$\|f([(x_{1},\pi(1)),\cdots,(x_{N},\pi(N))])-\mathcal{S}([(x_{1},\pi(1)),\cdots,(x_{N},\pi(N))])\|_{\infty}<\varepsilon$ (38)
for every $\{(x_{j},\pi(j))\}_{j\leq N}\in U$ and $\varepsilon>0$
Step 3. using DPOT to approximate the Sumformer.
In the following step, we will construct a DPOT to approximate the Sumformer $\mathcal{S}$ in a subset of $U$.
Lemma C.9.
Let $V=\{[(x_{1},1),\cdots,(x_{N},N)]|x_{i}\in\mathcal{X}\}\cong\mathcal{X}^{N}$ be a subset of $U$. Let $p:\mathcal{X}^{N}\to V$ be the projection mapping between two spaces. Then there exists a DPOT $\mathcal{N}$ that:
$\|\mathcal{N}(\{x_{i}\})-\mathcal{S}(p(\{x_{i}\}))\|_{\infty}<\varepsilon$ (39)
for every $\{x_{i}\}\in\mathcal{X}^{N}$ and $\varepsilon>0$
@@
In the upcoming proof, we will delve into the structure of DPOT and demonstrate how it can be employed to approximate the $\phi$ and $\psi$ within Sumformer.
Firstly, we will construct the lifting operator $\mathcal{R}$ as:
[…]
with the positional encoding $P(j)=\mathcal{F}^{-1}(\{1,\cdots,N\})(j)$. It is worth mentioning that, although we use the integer indices to replace to original vector indices, the discrete fourier transform is still defined on $\mathbb{T}^{d}$, not on a 1-dimensional sequence.
Then, in the first Fourier attention layer $\mathcal{L}_{1}$, we will first apply discrete Fourier Transform on the sequence, yielding:
$\begin{bmatrix}\frac{1}{\sqrt{N}}\sum_{k=1}^{N}\phi(x_{k})&\mathbf{0}_{d_{i}}&1\\ \cdots\\ \frac{1}{\sqrt{N}}\sum_{k=1}^{N}\phi(x_{k})e^{\frac{2\pi i\langle\mathbf{1},\mathbf{k}\rangle}{n}}&\mathbf{0}_{d_{i}}&N\end{bmatrix}\in\mathbb{R}^{N\times(2d_{i}+1)}$ (41)
After the Fourier transform, we will apply a pointwise 2-Layer MLP $\mathcal{K}_{1}$. According to the Universal Approximation Theorem of MLPs, we can find a $K_{1}$ that sufficiently fits the following function:
$K_{1}(\begin{bmatrix}\mathbf{z}&k\end{bmatrix})=\begin{cases}\begin{bmatrix}\mathbf{z}&k\end{bmatrix}&\text{if }k=1\\ \begin{bmatrix}0&k\end{bmatrix}&\text{if }k\neq 1\\ \end{cases}$ (42)
for any $\mathbf{z}\in\mathbb{R}^{2d_{i}}$.
Thus, the sequence will be mapped to
$K_{1}(\mathcal{R}(\{x_{k}\}))=\begin{bmatrix}\frac{1}{\sqrt{N}}\sum_{k=1}^{N}\phi(x_{k})&\mathbf{0}_{d_{i}}&1\\ \cdots\\ \mathbf{0}_{d_{i}}&\mathbf{0}_{d_{i}}&N\end{bmatrix}\in\mathbb{R}^{N\times(2d_{i}+1)}$ (43)
by $K_{1}$, and be transformed to
$\mathcal{F}^{-1}(K_{1}(\mathcal{R}(\{x_{k}\})))=\begin{bmatrix}\frac{1}{N}\Sigma&\mathbf{0}_{d_{i}}&P(1)\\ \cdots\\ \frac{1}{N}\Sigma&\mathbf{0}_{d_{i}}&P(N)\end{bmatrix}\in\mathbb{R}^{N\times(2d_{i}+1)}$ (44)
after discrete inverse Fourier Transform. Here, $\Sigma:=\sum_{k=1}^{N}\phi(x_{k})$.
The following mapping $M_{1}$ can be constructed as:
$M_{1}(\begin{bmatrix}\frac{1}{N}\Sigma&\mathbf{0}_{d_{i}}&P(k)\end{bmatrix})=\begin{bmatrix}\mathbf{0}_{d_{i}}&\Sigma&P(k)\end{bmatrix}$ (45)
so after the skipping connection, the sequence will finally become:
$\begin{bmatrix}x_{1}&\Sigma&2P(1)\\ \cdots\\ x_{N}&\Sigma&2P(N)\end{bmatrix}\in\mathbb{R}^{N\times(2d_{i}+1)}$ (46)
In the next Fourier layer $\mathcal{L}_{2}$, we kept the kernel $K_{2}$ as the identity mapping, ensuring the (inverse) Fourier transform has no effect on the sequence. Then we utilize the following MLP $M_{2}$ to approximate $\psi$:
$M_{2}(\begin{bmatrix}\Sigma&x_{k}&2P(k)\end{bmatrix})=\psi(\Sigma,x_{k})$ (47)
which is the desired output of Sumformer.
Appendix D Limitations of the work.
This work, as an empirical and numerical study, is capable of solving challenging PDE problems with high precision. However, due to the intricate mathematical nature of neural networks and the relative scarcity of theoretical research, employing them in risk-sensitive scenarios may pose safety risks.